40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 15 11:38:31 2009, rob, Update, Locking, MC_F channel dead mcfdead.png
    Reply  Thu Jul 16 14:41:38 2009, rob, Update, Locking, MC_F channel dead IOO_ICS_0_15.pngIOO_ICS_15_32.png
       Reply  Thu Jul 16 14:54:05 2009, rob, Update, Locking, MC_F channel dead 
Message ID: 1747     Entry time: Wed Jul 15 11:38:31 2009     Reply to this: 1758
Author: rob 
Type: Update 
Category: Locking 
Subject: MC_F channel dead 

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

Attachment 1: mcfdead.png  15 kB  Uploaded Wed Jul 15 12:52:14 2009  | Hide | Hide all
mcfdead.png
ELOG V3.1.3-